OpenStudy (anonymous):

Factor completely: 2x^3 + x^2 - 18x - 9

OpenStudy (anonymous):

F(x) = 2x^3 + x^2 - 18x - 9 First, try with x =1, 2, and 3 to see if f(1) = 0 ; f(2) = 0, or f(3) = 0 f(3) = 0, then f(x) can be factored by (x -3). Next, divide f(x) by (x - 3) f(x) = (x -3)(2x^2 + 7x + 3) The trinomial can be factored as: (x + 3)(2x + 1) Finally: f(x) = (x - 3)(x + 3) (2x + 1)